# Cislunar Constellation Design for Space Situational Awareness with Time-Expanded Facility Location Problem

### v1

@@ 5.3 Heuristics for Neighborhood Facility Swap
The major limitation of the LR procedure is the lack of convergence guarantee; depending on, among other things, the choice of constraints that are relaxed and the construction of the heuristic lower bound solution, it is possible that the procedure is no longer able to reduce the optimality gap. For the problem in question, preliminary experiments with the greedy/full-factorial allocation heuristics alone have revealed that while the algorithm performs well in most cases, there are instances where it gets stuck with intuitively suboptimal choice of observer locations. We thus develop additional heuristics to mutate the
@@
Using $\mathcal{C}_{j}$, the neighborhood swap heuristic which seeks to improve an existing solution by exploring neighboring candidate facility slots is conceived. This involves creating a new decision vector $Y_{\mathrm{new}}$ by mutating one of the active slot in the existing best feasible $Y_{\mathrm{best}}$, denoted as $j_{\mathrm{out}}$, to one of its neighboring slot in $\mathcal{C}_{j_{\mathrm{out}}}$, and evaluating the resulting feasible allocations $X_{\mathrm{new}}$ and $\theta_{\mathrm{new}}$ using the heuristic feasible allocation algorithm 1. This is repeated for each candidate neighboring slots $\mathcal{C}_{j}$ of each active slot in $Y_{\mathrm{best}}$. This procedure is summarized in Algorithm 4.
@@
8:         $Z_{\mathrm{new}}\leftarrow$ eqn. (15a) $\triangleright$ Evaluate objective
@@
Figure 9: Example of (a) intra-orbit and (b) inter-orbit neighborhood, shown for a 2:1 resonant L2 Southern Halo Orbit
@@ 5.3.1 Intra-Orbit Neighborhood
The intra-orbit neighborhood is defined based on proximity in angular location along a given LPO. Specifically, let $\Xi^{\alpha}_{j}\subset\mathcal{J}$ denote all other facility indices that are on the same LPO as $j$; for a given $j$, its LPO is discretized into $b_{j}$ slots according to equation (3), then $\Xi^{\alpha}_{j}$ has $b_{j}-1$ elements, excluding $j$ itself. We further define $\Delta\varphi_{j\xi}$ as the angular separation between facility locations $j$ and $\xi\in{\Xi}^{\alpha}_{j}$. Then, we define the permutation of $\xi$, denoted by $\pi_{C^{\alpha}}$, for which the absolute value of $\Delta\phi_{j\xi}$ is in ascending order,
$|\Delta\varphi_{j\pi_{C^{\alpha}}(1)}|\leq|\Delta\varphi_{j\pi_{C^{\alpha}}(2)}|\leq\ldots\leq|\Delta\varphi_{j\pi_{C^{\alpha}}(b_{j}-1)}|$ (37)
to arrive at the intra-orbit neighborhood $\mathcal{C}^{\alpha}_{j}$ of size $c^{\alpha}\leq b_{j}-1$, given by
$\mathcal{C}_{j}^{\alpha}=\left\{\xi_{\kappa}\,\mid\,\kappa=\pi_{C^{\alpha}}(1),\ldots,\pi_{C^{\alpha}}(c^{\alpha})\right\}$ (38)
Note that $c^{\alpha}$ is a tuning parameter for the heuristics that dictates the angular separation to be considered for swapping; while a larger $c^{\alpha}$ results in a more thorough evaluation of the neighborhood, the size of $\mathcal{C}_{j}^{\alpha}$ directly impacts the number of iterations of the inner for-loop in algorithm 4, and must thus be kept reasonably low. It is also recommended to choose $c^{\alpha}$ as a multiple of 2 to evenly explore candidate locations $\xi$ that either lead or lag in angular location relative to the originally chosen $j$.
@@ 5.3.2 Inter-Orbit Neighborhood
To arrive at an expression for the inter-orbit neighborhood of $j$, we begin by defining by $\mathcal{P}_{j}$ the set of LPOs that have the same $M$:$N$ resonance as the LPO of $j$. Then, for each candidate facility location $\xi$ along each LPO $\rho\in\mathcal{P}_{j}$, we determine the slot $\xi^{*}_{\rho}$ that has the smallest angular separation with respect to a reference target at position $\bar{\boldsymbol{r}}_{k}$ and for a given the line-of-sight vector from the Sun to target $k$, $\boldsymbol{l}_{Sk}$,
@@
Note that the choice of $\boldsymbol{l}_{Sk}$ is a direct consequence of choosing a reference epoch. In this work, we choose the initial epoch with Earth-Moon-Sun alignment as the reference, and define the reference target position as the weighted sum of all considered targets across all times,
$\bar{\boldsymbol{r}}_{k}=\dfrac{1}{\ell q}\sum_{t=1}^{\ell}\sum_{k=1}^{q}\boldsymbol{r}_{k}(t)$ (40)
Finally, the inter-orbit neighborhood $\mathcal{C}_{j}$ is given by
$\mathcal{C}_{j}=\left\{\xi^{*}_{\rho}\,\mid\,\forall\rho\in\mathcal{P}_{j}\right\}$ (41)
Note that the inter-orbit neighborhood swap is only attempted if the LR and the corresponding feasible solution from algorithm 1 fail to reduce the optimality gap over a pre-defined number of iterations; this is due to the computational expense of this swap approach, requiring to run algorithm 1 for each candidate solution, as well as to avoid prematurely swapping to different LPO slots before the nearby slots in phase along a given LPO has been sufficiently explored.
5.4 Subgradient Optimization
For the dual problem, we seek to obtain multipliers $\lambda_{jt}$ and $\eta_{tk}$ that yields the tightest upper bound,
$Z_{\mathrm{LR}}^{*}=\min_{\lambda_{jt},\eta_{tk}}Z_{\mathrm{LR}}$ (42)
The Lagrange multipliers are updated using the subgradient method
@@ 5.4 Subgradient Optimization
where $s^{(h)}$ is the step-size given by
$s^{(h)}=\dfrac{\mu^{(h)}(Z_{\mathrm{LR}}-Z_{\mathrm{LH}})}{\left[\sum_{j=1}^{n}\sum_{t=1}^{\ell}\max\left(0,\sum_{i=1}^{m}M_{ijtk}\bar{X}_{ijt}-1\right)\right]^{2}\left[\sum_{t=1}^{\ell}\sum_{k=1}^{q}\max\left(0,\bar{\theta}_{tk}-\sum_{i=1}^{m}\sum_{j=1}^{n}M_{ijtk}\bar{X}_{ijt}\right)\right]^{2}}$ (45)
In equation (45), $\mu^{(h)}$ is a hyperparameter for scaling the step-size, which is reduced if no reduction of the gap between the lower-bound and upper-bound solutions is observed; initially, $\mu^{(0)}=2.0$ is typically chosen when employing the subgradient update, with a step reduction factor of 0.5.
5.5 Implementation Considerations
@@ 5.5.2 Parallelism
For a given choice of facilities $Y$, the feasible allocation construction algorithm 1 makes allocation decisions $X_{ijt}$ for each time step, each of which is independent of allocation decisions at other time steps. This makes the algorithm well-suited for parallelism; we opt for the multiple-threading paradigm through Julia’s Threads.@threads decorator to the for-loop in line 3 through 20, as the operation within each loop is not too expensive, and we gain more benefit from a shared memory. Note that both the intra- and inter-orbit neighborhood swap schemes require recomputing the best heuristic allocations $X_{ijt}$ for each new hypothesized $Y$, thus a speed-up of algorithm 1 is particularly beneficial to speed up the overall Lagrangean method.
@@ 5.5.3 Memoization
During the heuristic process of the Lagrangean method, it is possible that the Lagrangean relaxed solution $\bar{Y}$ or the result of either an intra- and inter-orbit neighborhood swap result in a choice of facilities that has already been considered by the algorithm. We make use of function memoization on the feasible allocation construction algorithm 1 by storing the outputs $\underline{X}$ and $\underline{\theta}$ that results from a given set of active facility indices $\mathcal{J}^{\bar{Y}}$.
@@ 6 Numerical Results
The TE-$p$-MP is solved for various choices of observation demand as well as observer parameters. Each instance is solved using the LM and Gurobi, in order to study the efficacy of the proposed former approach. This Section initially focuses on comparing the performance of the two methods, emphasizing on the quality of solutions they yield. Then, closer attention is drawn to the obtained architectures for a select number of instances, in order to provide insights on the CSSA problem.
The three sets of demand presented in Section 3 are considered; the corresponding problem dimensions are summarized in Table 3. For each set of demand introduced in Section 3, we consider LPOs given in Table 9, with an approximate time-spacing parameter $\Delta t_{b}=12$ hours. Using equation (3), this results in $n=1212$ candidate facility slots in total, summarized in Table 9. For the system with all observers to return to their original position, the considered set of resonances requires 4 synodic months. A time-step of 60 steps per synodic month, corresponding to roughly 24 hours, is selected, thus resulting in $\ell=120$ time-steps. All problem instances are solved on a desktop PC running on eight i7-10700 CPU cores at 2.90GHz. Table 4 gives the hyperparameters for the LM; Table 5 gives the non-default options set with Gurobi version 9.5.0.
@@
Demand SOI Cone of shame LET transit
Number of pointing directions $m$ 14 14 14
Number of facility locations $n$ 1212 1212 1212
Number of time steps $\ell$ 120 120 120
Number of targets $q$ 120 304 675
Fraction of entries in $M^{b}$ $\bar{m}_{\mathrm{crit}}=15$ 0.0111 0.0047 0.0007
FOV $60^{\circ}$ $\bar{m}_{\mathrm{crit}}=18$ 0.0496 0.0460 0.0043
$\bar{m}_{\mathrm{crit}}=20$ 0.0572 0.0581 0.0049
$\bar{m}_{\mathrm{crit}}=15$ 0.0421 0.0180 0.0026
FOV = $120^{\circ}$ $\bar{m}_{\mathrm{crit}}=18$ 0.1881 0.1711 0.0176
$\bar{m}_{\mathrm{crit}}=20$ 0.2198 0.2177 0.0204
6.1 Comparison of Lagrangean Method against Gurobi
@@
We first compare the proposed Lagrangean method against Gurobi, for TE-$p$-MP instances with static and dynamic demands. For the sake of a fair comparison, we set a time limit of 500 seconds on both approaches to assess their performance. In addition, the solution obtained by running Gurobi with a time limit of 3600 seconds is reported. Note that while extending the time limit may allow Gurobi to find better solutions, this is not necessarily desirable, especially when various problem parameters, such as the number of satellites $p$, cut of magnitude, or FOV, must be explored and optimized. Nevertheless, the solution from Gurobi with a 3600 seconds time limit serves as a proxy of a near-optimal solution obtained with a reasonably long time limit in most instances.
@@
Tables 6(c)7(c), and 8(c) report the sum of targets observed across all time steps, $\Theta$, and the actual CPU solve time for each of the three observation demands, respectively. In all three cases, the solutions obtained with Gurobi using time limits of 500 and 3600 seconds and with the LM using a time limit of 500 seconds are reported. In addition, the percentage difference of the Lagrangean method solution with respect to each Gurobi solution is reported. Note that the actual CPU solve time may exceed 500 seconds for both Gurobi and the Lagrangean method; in Gurobi, this is due to the computation of attributes associated with the terminated solution, while with the Lagrangean method, the solve time is checked between the generation of feasible lower bound solution, which may start before the solve time limit, but exceed it before another solve time check happens.
@@
Across the three demands, we first note that LM finds a superior solution to Gurobi with the 500 seconds time limit (i.e., under a fair condition) in all cases with $\bar{m}_{\mathrm{crit}}=20$, and in all except two cases with $\bar{m}_{\mathrm{crit}}=18$, the latter of which corresponding to $p=2$ with the Cone of Shame and LET demands. For cases with $\bar{m}_{\mathrm{crit}}=15$, Gurobi finds competitive solutions for the SOI and Cone of Shame demands, while it is found to struggle with the LET demand for $p\geq 3$. Letting Gurobi run for up to 3600 seconds improves the solution in some instances, but the behavior is hard to predict. For example, in the SOI demand, cases with $\bar{m}_{\mathrm{crit}}=18$ and $p=3$ has a solution that improves from $\Theta_{500}=0.2802$ to $\Theta_{3600}=0.9885$, while with $\bar{m}_{\mathrm{crit}}=18$ and $p=4$, the solution remains at $\Theta_{500}=\Theta_{3600}=0.2863$. The performance of Gurobi is largely dictated by how effective the underlying B&B algorithm can prune the nodes as well as the presolve algorithm, both of which are in turn impacted by the solver’s hyperparameters.
@@
The performance of Gurobi, both after 500 and 3600 seconds, may thus be seen as a measure of the difficulty of the TE-$p$-MP instance; the problem appears to be easy when the observation is difficult, with observers are few (low $p$) and/or are equipped with worse sensors (lower $\bar{m}_{\mathrm{crit}}$), or when the observation is easy, with many observers (high $p$) and/or high-performance sensors (higher $\bar{m}_{\mathrm{crit}}$), resulting in an instance where the observation task is “easy”. Examples of the former include the cases with $\bar{m}_{\rm crit}=15,18$ and $p=2,3$ for the LET demand, while examples of the latter include the cases with $\bar{m}_{\rm crit}=15$ and $p=4,5$, also for the LET demand. Instances of intermediate observation difficulty, such as the cases with $\bar{m}_{\rm crit}=18,20$ and $p=2\sim 5$ for the Cone of Shame demand, are found to result in the poorest performances from Gurobi.
@@
In contrast, the performance of the LM is found to be consistent in terms of observation condition: increasing $p$ always increases $\Theta_{\rm LH}$, while increasing $\bar{m}_{\mathrm{crit}}$ also increases $\Theta_{\rm LH}$ except for the two cases with the Cone of Shame demand with $p=2,3$ going from $\bar{m}_{\mathrm{crit}}=18$ to $20$, albeit with a small reduction, from $0.7334$ to $0.7297$ with $p=2$, and from $0.8192$ to $0.8080$ with $p=3$. Note that these two instances correspond to some of the cases where Gurobi struggles the most. Among the 36 instances, 11 instances resulted in an LM solution that was inferior to Gurobi after 500 seconds; within them, only 4 cases resulted in a solution that was worse by over 10%. Overall, the consistency of the LM offers an attractive reliable approach for assessing the achievable observation from a given TE-$p$-MP instance. Alternatively, one may consider using Gurobi with a relatively short solve time limit such as 500 seconds, and use the LM as a consistent fallback solver to recover descent solutions in case Gurobi under-performs.
Table 4: Lagrangean method hyperparameters
@@
Table 5: Non-default Gurobi options
Option Value
Gap tolerance 0.01
Method 2
MIPFocus 2
Table 6: TE-$p$-MP instance with SOI demand and for FOV = $60^{\circ}$ solved using Gurobi (MIPGap 1%) versus Lagrangean method. All reported times are in seconds.
@@
$p$ Gurobi (TimeLimit $=500$) Gurobi (TimeLimit $=3600$) Lagrangean Method
@@
2 0.1558 316 OPTIMAL - - - 0.1506 507 $-3.34$ $-3.34$
3 0.2238 500 TIME_LIMIT 0.2238 3600 TIME_LIMIT 0.2057 553 $-8.10$ $-8.10$
4 0.2653 500 TIME_LIMIT 0.2741 3601 TIME_LIMIT 0.2600 543 $-2.02$ $-5.14$
5 0.3131 500 TIME_LIMIT 0.3135 3600 TIME_LIMIT 0.3104 551 $-0.86$ $-1.00$
@@
$p$ Gurobi (TimeLimit $=500$) Gurobi (TimeLimit $=3600$) Lagrangean Method
@@
2 0.8567 503 TIME_LIMIT 0.9160 3601 TIME_LIMIT 0.5506 503 $-35.72$ $-39.89$
3 0.2802 500 TIME_LIMIT 0.9885 3600 TIME_LIMIT 0.7007 515 $+150.06$ $-29.11$
4 0.2863 500 TIME_LIMIT 0.2863 3600 TIME_LIMIT 0.7858 534 $+174.46$ $+174.46$
5 0.3665 500 TIME_LIMIT 0.3665 3600 TIME_LIMIT 0.9383 534 $+156.04$ $+156.04$
@@
$p$ Gurobi (TimeLimit $=500$) Gurobi (TimeLimit $=3600$) Lagrangean Method
@@
2 0.4176 503 TIME_LIMIT 0.9367 3721 TIME_LIMIT 0.5791 573 $+38.66$ $-38.18$
3 0.4383 504 TIME_LIMIT 0.9662 3603 TIME_LIMIT 0.6514 564 $+48.61$ $-32.58$
4 0.4690 500 TIME_LIMIT 0.9173 3601 TIME_LIMIT 0.7949 533 $+69.47$ $-13.35$
5 0.5332 500 TIME_LIMIT 0.9926 1633 OPTIMAL 0.8743 588 $+63.97$ $-11.92$
Table 7: TE-$p$-MP instance with Cone of Shame demand and FOV = $60^{\circ}$ solved using Gurobi (MIPGap 1%) versus Lagrangean method. All reported times are in seconds.

### v2

@@ 5.3 Heuristics for Neighborhood Facility Swap
The major limitation of the relaxation procedure is the lack of convergence guarantee; depending on, among other things, the choice of constraints that are relaxed and the construction of the heuristic lower bound solution, it is possible that the procedure is no longer able to reduce the optimality gap. For the problem in question, preliminary experiments with the greedy/full-factorial allocation heuristics alone have revealed that while the algorithm performs well in most cases, there are instances where it gets stuck with intuitively suboptimal choice of observer locations. We thus develop additional heuristics to mutate the chosen set of facilities.
@@
Using $\mathcal{C}_{j}$, the neighborhood swap heuristic which seeks to improve an existing solution by exploring neighboring candidate facility slots is conceived. The neighborhood swap involves creating a new decision vector $Y_{\mathrm{new}}$ by mutating one of the active slot in the existing best feasible $Y_{\mathrm{best}}$, denoted by $j_{\mathrm{out}}$, to one of its neighboring slot in $\mathcal{C}_{j_{\mathrm{out}}}$, and evaluating the resulting feasible allocations $X_{\mathrm{new}}$ and $\theta_{\mathrm{new}}$ using the heuristic feasible allocation Algorithm 2. This process is repeated for each candidate neighboring slots $\mathcal{C}_{j}$ of each active slot in $Y_{\mathrm{best}}$. The neighborhood swap heuristic procedure is summarized in Algorithm 5.
@@
8:         $Z_{\mathrm{new}}\leftarrow$ eqn. (24a) $\triangleright$ Evaluate linear objective function
@@
Figure 9: Example of intra- and inter-orbit neighborhoods of a slot on the 2:1 resonant L2 Southern Halo Orbit
@@ 5.3.1 Intra-Orbit Neighborhood
The intra-orbit neighborhood is defined based on proximity in angular location along a given LPO. Specifically, let $\Xi_{j}\subset\mathcal{J}$ denote all other facility indices that are on the same LPO as $j$; for a given $j$, its LPO is discretized into $b_{j}$ slots according to equation (13), then $\Xi_{j}$ has $b_{j}-1$ elements, excluding $j$ itself. We further define $\Delta\varphi_{j\xi}$ as the angular separation between facility locations $j$ and $\xi\in\Xi_{j}$. Then, we define the permutation of $\xi$, denoted by $\pi_{\rm intra}$, for which the absolute value of $\Delta\phi_{j\xi}$ is in ascending order,
$|\Delta\varphi_{j\pi_{\rm intra}(1)}|\leq|\Delta\varphi_{j\pi_{\rm intra}(2)}|\leq\ldots\leq|\Delta\varphi_{j\pi_{\rm intra}(b_{j}-1)}|$ (46)
We finally construct the intra-orbit neighborhood $\mathcal{C}^{\rm intra}_{j}$ of size $c^{\rm intra}\leq b_{j}-1$ by taking the first $c^{\rm intra}$ elements in the permutation $\pi_{\rm intra}$,
$\mathcal{C}_{j}^{\rm intra}=\left\{\xi_{\kappa}\,\mid\,\kappa=\pi_{\rm intra}(1),\ldots,\pi_{\rm intra}(c^{\rm intra})\right\}$ (47)
Note that $c^{\rm intra}$ is a tuning parameter for the heuristics that dictates the angular separation to be considered for swapping; while a larger $c^{\rm intra}$ results in a more thorough evaluation of the neighborhood, the size of $\mathcal{C}_{j}^{\rm intra}$ directly impacts the number of iterations of the inner for-loop in Algorithm 5, and must thus be kept reasonably low. It is also recommended to choose $c^{\rm intra}$ as a multiple of 2 to evenly explore candidate locations $\xi$ that either lead or lag in angular location relative to the originally chosen $j$.
@@ 5.3.2 Inter-Orbit Neighborhood
To arrive at an expression for the inter-orbit neighborhood of $j$, we begin by defining by $\mathcal{P}_{j}$ the set of LPOs that have the same $M$:$N$ resonance as the LPO of $j$. We also define a reference target position $\boldsymbol{r}_{k_{\rm ref}}$; in this work, we choose the mean location across all targets, $\boldsymbol{r}_{k_{\rm ref}}=\sum_{k}\boldsymbol{r}_{k}$. Then, for each candidate facility location $\xi$ along each LPO $\rho\in\mathcal{P}_{j}$, we determine the slot $\xi^{*}_{\rho}$ that achieves the closest solar phase angle with respect to $\boldsymbol{r}_{k_{\rm ref}}$ as the original slot $j$, given by
@@
The choice of $\boldsymbol{l}_{Sk}$ is a direct consequence of choosing a reference epoch. In this work, we choose the initial epoch with Earth-Moon-Sun alignment as the reference, and define the reference target position as the weighted sum of all considered targets across all times,
$\boldsymbol{r}_{k_{\rm ref}}=\dfrac{1}{\ell q}\sum_{t=1}^{\ell}\sum_{k=1}^{q}\boldsymbol{r}_{k}(t)$ (49)
Finally, the inter-orbit neighborhood $\mathcal{C}_{j}^{\rm inter}$ is given by
$\mathcal{C}_{j}^{\rm inter}=\left\{\xi^{*}_{\rho}\,\mid\,\forall\rho\in\mathcal{P}_{j}\right\}$ (50)
The inter-orbit neighborhood swap is only attempted if the relaxed solution and the corresponding feasible solution from Algorithm 2 fail to reduce the optimality gap over a pre-defined number of iterations. We place this restriction due to two considerations: firstly, we avoid prematurely swapping to different LPO slots before the nearby slots in phase along a given LPO has been sufficiently explored. Secondly, the inter-orbit neighborhood Algorithm 5 requires running the intra-orbit neighborhood Algorithm 2 for each candidate solution, thus making it more computationally intensive.
5.4 Lagrange Multipliers Update with Subgradient Method
The Lagrange multipliers are updated according to the subgradient method according to (8); for the TE-$p$-MP, the multiplier update formulae are
@@ 5.4 Lagrange Multipliers Update with Subgradient Method
where the expression for the step-size (9) becomes
$s^{(h)}=\dfrac{\mu^{(h)}(Z_{\mathrm{relax}}-Z_{\mathrm{feas}})}{\left[\sum_{j=1}^{n}\sum_{t=1}^{\ell}\max\left(0,\sum_{i=1}^{m}M_{ijtk}\bar{X}_{ijt}-1\right)\right]^{2}\left[\sum_{t=1}^{\ell}\sum_{k=1}^{q}\max\left(0,\bar{\theta}_{tk}-\sum_{i=1}^{m}\sum_{j=1}^{n}M_{ijtk}\bar{X}_{ijt}\right)\right]^{2}}$ (53)
5.5 Considerations for Implementing the Lagrangian Method
@@ 5.5.2 Parallelism
For a given choice of facilities $Y$, the feasible allocation construction Algorithm 2 makes allocation decisions $X_{ijt}$ for each time step, each of which is independent of allocation decisions at other time steps. This makes the algorithm well-suited for parallelism; we opt for the multiple-threading paradigm through Julia’s Threads.@threads decorator to the for-loop in line 3 through 20, as the operation within each loop is not too expensive, and we gain more benefit from a shared memory. Note that both the intra- and inter-orbit neighborhood swap schemes require recomputing the best heuristic allocations $X_{ijt}$ for each new hypothesized $Y$, thus a speed-up of Algorithm 2 is particularly beneficial to speed up the overall Lagrangian method.
@@ 5.5.3 Memoization
During the heuristic process of the Lagrangian method, it is possible that the Lagrangian relaxed solution $\bar{Y}$ or the result of either an intra- and inter-orbit neighborhood swap result in a choice of facilities that has already been considered by the algorithm. We make use of function memoization on the feasible allocation construction Algorithm 2 by storing the outputs $\underline{X}$ and $\underline{\theta}$ that results from a given set of active facility indices $\mathcal{J}^{\bar{Y}}$.
@@ 6 Numerical Results
The TE-$p$-MP is solved for various choices of observation demand as well as observer parameters. Each instance is solved using the LM and B&B via Gurobi. Gurobi is a state-of-the-art optimization suite with a particular focus on B&B mixed-integer linear programs and serves as a point of comparison to study the efficacy of the proposed LM. This Section initially focuses on comparing the performance of the two methods, emphasizing on the quality of solutions they yield. Then, closer attention is drawn to the obtained architectures for a select number of instances, in order to provide insights on the CSSA problem.
The two sets of demand presented in Section 3 are considered. The corresponding problem dimensions are summarized in Table 3. For each set of demand introduced in Section 3, we consider LPOs given in Table 7, with an approximate time-spacing parameter $\Delta t_{b}=12$ hours. Using equation (13), this results in $n=1212$ candidate facility slots in total, summarized in Table 7. Since the set of candidate observer LPOs is assumed to be resonant, any architecture involving any combination of LPOs will return to its original configuration after $N_{\rm syn}$ synodic period, as introduced in Section 2.3.2. Leveraging this feature, we consider a TE-$p$-MP with a time horizon of $N_{\rm syn}=4$ synodic months. Each synodic month is discretized with a time-step of 30 steps per synodic month, corresponding to roughly 24 hours, thus resulting in $\ell=120$ time-steps. Note that the $N_{\rm syn}$-synodic period-repeating feature of the candidate LPOs is not necessitated by the TE-$p$-MP formulation itself; one may choose to formulate the problem with non-repeating candidate observer locations, and set a longer time-horizon as well.
All problem instances are solved on a desktop PC running on 20 i7-12700 CPU cores at 2.10GHz base speed. Table 4 gives the hyperparameters for the LM. For B&B, we employ Gurobi version 11.0.1, using the barrier method to solve the initial root relaxation as it was found through preliminary experiments to improve the algorithm’s performance to TE-$p$-MP instances. The optimality gap tolerance is set to $1\%$. All other hyperparameters are left unchanged from their default values.
@@
Demand Cone of shame LET transit
Number of pointing directions $m$ 14 14
Number of facility locations $n$ 1212 1212
Number of time steps $\ell$ 120 120
Number of targets $q$ 304 675
Fraction of entries in $M^{b}$ $\bar{m}_{\mathrm{crit}}=15$ 0.0047 0.0007
FOV $=60^{\circ}$ $\bar{m}_{\mathrm{crit}}=18$ 0.0460 0.0043
$\bar{m}_{\mathrm{crit}}=20$ 0.0581 0.0049
$\bar{m}_{\mathrm{crit}}=15$ 0.0180 0.0026
FOV $=120^{\circ}$ $\bar{m}_{\mathrm{crit}}=18$ 0.1711 0.0176
$\bar{m}_{\mathrm{crit}}=20$ 0.2177 0.0204
6.1 Comparison of Lagrangian Method against Branch-and-Bound
@@
We first compare the proposed Lagrangian method against Gurobi’s B&B, for TE-$p$-MP instances with static and dynamic demands. For the sake of a fair comparison, we set a time limit of 500 seconds on both approaches to assess their performance. In addition, the solution obtained by Gurobi with a time limit of 3600 seconds is reported. Note that while extending the time limit may allow B&B to find better solutions, this is not necessarily desirable, especially when various problem parameters, such as the number of satellites $p$, cut of magnitude, or FOV, must be explored and optimized. Nevertheless, the solution from Gurobi with a 3600-second time limit serves as a proxy of a near-optimal solution obtained with a reasonably long time limit in most instances.
@@
Tables 5(c), and 6(c) report the sum of targets observed across all time steps, $\Theta$, and the actual CPU solve time for each of the two observation demands, respectively. In both cases, the solutions obtained with Gurobi using time limits of 500 and 3600 seconds and with the LM using a time limit of 500 seconds are reported. In addition, the percentage difference of the Lagrangian method solution with respect to each B&B solution is reported. Note that the actual CPU solve time may exceed 500 seconds for both Gurobi and the Lagrangian method; in Gurobi, this is due to the computation of attributes associated with the terminated solution, while with the Lagrangian method, the solve time is checked between the generation of feasible lower bound solution, which may start before the solve time limit, but exceed it before another solve time check happens.
@@
Across both demands, we first note that LM finds a superior solution to B&B with the 500 seconds time limit (i.e., under a fair condition) in all cases with $\bar{m}_{\mathrm{crit}}=20$, and in all except two cases with $\bar{m}_{\mathrm{crit}}=18$, the latter of which corresponding to $p=2$ with the Cone of Shame and LET demands. For cases with $\bar{m}_{\mathrm{crit}}=15$, B&B finds competitive solutions for the Cone of Shame demand, while it is found to struggle with the LET demand for $p\geq 3$. Letting Gurobi run for up to 3600 seconds improves the solution in some instances, but the behavior is hard to predict. For example, in the Cone of Shame demand, cases with $\bar{m}_{\mathrm{crit}}=18$ and $p=3$ has a solution that improves from $\Theta_{500}=0$ to $\Theta_{3600}=0.6355$, while with $\bar{m}_{\mathrm{crit}}=18$ and $p=4$, the solution improves only from $\Theta_{500}=0.6556$ to $\Theta_{3600}=0.6975$.
@@
The performance of B&B, both after 500 and 3600 seconds, may be seen as a measure of the difficulty of the TE-$p$-MP instance. The problem appears to be easy when the overall visibility of targets is low, corresponding to instances with fewer (low $p$) observers and/or are equipped with worse sensors (lower $\bar{m}_{\mathrm{crit}}$), or when the observation is easy, with many observers (high $p$) and/or high-performance sensors (higher $\bar{m}_{\mathrm{crit}}$), resulting in an instance where the observation task is “easy”. Examples of the former include the cases with $\bar{m}_{\rm crit}=15,18$ and $p=2,3$ for the LET demand, while examples of the latter include the cases with $\bar{m}_{\rm crit}=15$ and $p=4,5$, also for the LET demand. Instances of intermediate observation difficulty, such as the cases with $\bar{m}_{\rm crit}=18,20$ and $p=2,3,4,5$ for the Cone of Shame demand, are found to result in the poorest performances from B&B.
@@
The performance of the LM is consistent in terms of observation condition: increasing $p$ always increases $\Theta_{\rm LH}$, while increasing $\bar{m}_{\mathrm{crit}}$ also increases $\Theta_{\rm LH}$ except for the two cases with the Cone of Shame demand with $p=2,3$ going from $\bar{m}_{\mathrm{crit}}=18$ to $20$, albeit with a small reduction, from $0.7334$ to $0.7297$ with $p=2$, and from $0.8192$ to $0.8080$ with $p=3$. Note that these two instances correspond to some of the cases where B&B struggles the most. Among the 24 instances, 6 instances resulted in an LM solution that was inferior to B&B after the 500-second time limit; within them, only 4 cases resulted in a solution that was worse by over 10%. Overall, the consistency of the LM offers an attractive reliable approach for assessing the achievable observation from a given TE-$p$-MP instance. Alternatively, one may consider using Gurobi with a relatively short solve time limit such as 500 seconds, and use the LM as a consistent fallback solver to recover decent solutions in case B&B under-performs.
Table 4: Lagrangian method hyperparameters
@@
Table 5: TE-$p$-MP instance with Cone of Shame demand and FOV = $60^{\circ}$ solved using Branch-and-Bound versus Lagrangian method. All reported times are in seconds.
@@
$p$ B&B (TimeLimit $=500$) B&B (TimeLimit $=3600$) Lagrangian Method
@@
2 0.0419 101 OPTIMAL - - - 0.0402 620 $-3.99$ $-3.99$
3 0.0608 501 TIME_LIMIT 0.0608 681 OPTIMAL 0.0533 575 $-12.22$ $-12.22$
4 0.0778 500 TIME_LIMIT 0.0783 3601 TIME_LIMIT 0.0717 614 $-7.89$ $-8.54$
5 0.0939 500 TIME_LIMIT 0.0939 3601 TIME_LIMIT 0.0824 603 $-12.29$ $-12.29$
@@
$p$ B&B (TimeLimit $=500$) B&B (TimeLimit $=3600$) Lagrangian Method
@@
2 0.6058 501 TIME_LIMIT 0.7259 3601 TIME_LIMIT 0.6341 542 $+4.67$ $-12.65$
3 0.0000 501 TIME_LIMIT 0.6355 3601 TIME_LIMIT 0.7786 584 $+Inf$ $+22.52$
4 0.6756 501 TIME_LIMIT 0.6975 3601 TIME_LIMIT 0.8670 634 $+28.32$ $+24.31$
5 0.0000 500 TIME_LIMIT 0.7576 3602 TIME_LIMIT 0.9142 703 $+Inf$ $+20.68$
@@
$p$ B&B (TimeLimit $=500$) B&B (TimeLimit $=3600$) Lagrangian Method
@@
2 0.0000 501 TIME_LIMIT 0.6180 3602 TIME_LIMIT 0.6328 536 $+Inf$ $+2.39$
3 0.0000 501 TIME_LIMIT 0.8577 3601 TIME_LIMIT 0.7223 593 $+Inf$ $-15.79$
4 0.0000 501 TIME_LIMIT 0.7854 3602 TIME_LIMIT 0.9271 597 $+Inf$ $+18.04$
5 0.0000 501 TIME_LIMIT 0.9558 3603 TIME_LIMIT 0.9599 584 $+Inf$ $+0.43$
Table 6: TE-$p$-MP instance with LET demand and FOV = $60^{\circ}$ solved using Branch-and-Bound versus Lagrangian method. All reported times are in seconds.